# Mean-Field Sampling for Cooperative Multi-Agent Reinforcement Learning

### v1

@@ 2.2 Technical Background
The solution offered by mean-field value iteration and standard $Q$-learning requires a sample complexity of $\min\{\widetilde{O}(|\mathcal{S}_{g}||\mathcal{A}_{g}||\mathcal{Z}_{l}|^{n}),\widetilde{O}(|\mathcal{S}_{g}||\mathcal{A}_{g}||\mathcal{Z}_{l}|n^{|\mathcal{Z}_{l}|})\}$, where one uses standard $Q$-learning when $|\mathcal{Z}_{l}|^{n-1}<n^{|\mathcal{Z}_{l}|}$, and mean-field value iteration otherwise. In each of these regimes, as $n$ scales, the update complexity can become incredibly computationally intensive. Therefore, we introduce the SUBSAMPLE-MFQ algorithm to mitigate the cost of scaling the number of local agents.
3 Method and Theoretical Results
In this section, we propose the SUBSAMPLE-MFQ algorithm to overcome the polynomial (in $n$) sample complexity of mean-field value iteration and the exponential (in $n$) sample complexity of traditional $Q$-learning. In our algorithm, the global agent randomly samples a subset of local agents $\Delta\subseteq[n]$ such that $|\Delta|=k$, for $k\leq n$. It ignores all other local agents $[n]\setminus\Delta$, and performs value iteration to learn the $Q$-function $\widehat{Q}_{k,m}^{*}$ and policy $\widehat{\pi}_{k,m}^{*}$ for this surrogate subsystem of $k$ local agents, where $m$ is the number of samples used to update the $Q$-functions’ estimates of the unknown system. Here, When $|\mathcal{Z}_{l}|^{k-1}<k^{|\mathcal{Z}_{l}|}$, the algorithm uses traditional value-iteration (Algorithm 1), and when $|\mathcal{Z}_{l}|^{k-1}>k^{|\mathcal{Z}_{l}|}$, it uses mean-field value iteration (Algorithm 2). The surrogate reward gained by this subsystem at each time step is $r_{\Delta}:\mathcal{S}\times\mathcal{A}\to\mathbb{R}$:
To convert the optimality of each agent’s action within the $k$ local-agent subsystem to an approximate optimality on the full $n$-agent system, we use a randomized policy ${\pi}^{\mathrm{est}}_{k,m}$ (Algorithm 3), where the global agent samples $\Delta\in\mathcal{U}\binom{[n]}{k}$ at each time-step to derive the action $a_{g}\leftarrow\widehat{\pi}_{k,m}^{*}(s_{g},s_{\Delta})$, and where each local $i$ agent samples $k-1$ other local agents $\Delta_{i}$ to derive the action $\widehat{\pi}_{k,m}^{*}(s_{g},s_{i},s_{\Delta_{i}})$. Finally, Theorem 3.4 shows that the policy ${\pi}_{k,m}^{\mathrm{est}}$ converges to the optimal policy $\pi^{*}$ as $k\to n$. We first present Algorithms 2 and 1 (SUBSAMPLE-MFQ: Learning) and Algorithm 3 (SUBSAMPLE-MFQ: Execution), which we describe below. For this, a crucial characterization is the notion of the empirical distribution function:
Definition 3.1 (Empirical Distribution Function).
For any population $(z_{1},\dots,z_{n})\in\mathcal{Z}_{l}^{n}$, where $\mathcal{Z}_{l}\coloneq\mathcal{S}_{l}\times\mathcal{A}_{l}$, define the empirical distribution function $F_{z_{\Delta}}:\mathcal{Z}_{l}\to\mathbb{R}_{+}$ for $\Delta\subseteq[n]$ such that $|\Delta|=k$ by:
$F_{z_{\Delta}}(x)\coloneq F_{s_{\Delta},a_{\Delta}}(x)\coloneq\frac{1}{k}\sum_{i\in\Delta}\mathbbm{1}\{s_{i}=z_{s},a_{i}=z_{a}\},\quad\forall z\coloneq(z_{s},z_{a})\in\mathcal{S}_{l}\times\mathcal{A}_{l}.$ (8)
Let $\mu_{k}(\mathcal{Z}_{l})\!\coloneq\!\left\{\frac{b}{k}|b\in\{0,\dots,k\}\right\}^{|\mathcal{S}_{l}|\times|\mathcal{A}_{l}|}$ be the space of $|\mathcal{S}_{l}|\times|\mathcal{A}_{l}|$-length vectors where each entry in a vector is an element of $\{0,\frac{1}{k},\frac{2}{k},\dots,1\}$ such that $F_{z_{\Delta}}\in\mu_{k}(\mathcal{Z}_{l})$. Here, $F_{z_{\Delta}}$ is the proportion of agents in the $k$-local-agent subsystem at each state.
Algorithms 2 and 1 (Offline learning). Let $m\in\mathbb{N}$ denote the sample size for the learning algorithm with sampling parameter $k\leq n$. When $|\mathcal{Z}_{l}|^{k-1}\!\leq\!k^{|\mathcal{Z}_{l}|}$, we empirically learn the optimal $Q$-function for a subsystem with $k$-local agents denoted by $\widehat{Q}_{k,m}^{\mathrm{est}}\!:\!\mathcal{S}_{g}\times\mathcal{S}_{l}^{k}\times\mathcal{A}_{g}\times\mathcal{A}_{l}^{k}\to\mathbb{R}$: set $\widehat{Q}_{k,m}^{0}(s_{g},s_{\Delta},a_{g},a_{\Delta})\!=\!0$ for all $(s_{g},s_{\Delta},a_{g},a_{\Delta})\!\in\!\mathcal{S}_{g}\times\mathcal{S}_{l}^{k}\times\mathcal{A}_{g}\times\mathcal{A}_{l}^{k}$. At time step $t$, set $\widehat{Q}_{k,m}^{t+1}(s_{g},{s_{\Delta}},a_{g},a_{\Delta})={\widetilde{\mathcal{T}}}_{k,m}\widehat{Q}_{k,m}^{t}(s_{g},s_{\Delta},a_{g},a_{\Delta})$, where ${\widetilde{\mathcal{T}}}_{k,m}$ is the empirically adapted Bellman operator in Eq. 9.
When $|\mathcal{Z}_{l}|^{k-1}\!>\!k^{|\mathcal{Z}_{l}|}$, we empirically learn the optimal mean-field $Q$-function for a $k$ local agent system, denoted (with abuse of notation) by $\widehat{Q}_{k,m}^{\mathrm{est}}\!:\!\mathcal{S}_{g}\!\times\!\mathcal{S}_{1}\times\mu_{k-1}(\mathcal{Z}_{l})\times\mathcal{A}_{g}\times\mathcal{A}_{l}\to\mathbb{R}$. For all $(s_{g},s_{1},F_{z_{\widetilde{\Delta}}},a_{1},a_{g})\in\mathcal{S}_{g}\times\mathcal{S}_{l}\times\mu_{k-1}(\mathcal{Z}_{l})\times\mathcal{A}_{l}\times\mathcal{A}_{g}$, set $\widehat{Q}_{k,m}^{0}(s_{g},s_{1},F_{z_{\widetilde{\Delta}}},a_{1},a_{g})\!=\!0$. At time $t$, set $\widehat{Q}_{k,m}^{t+1}(s_{g},s_{1},F_{z_{\widetilde{\Delta}}},a_{1},a_{g})\!=\!\widehat{\mathcal{T}}_{k,m}\widehat{Q}_{k,m}^{t}(s_{g},s_{1},F_{z_{\widetilde{\Delta}}},a_{1},a_{g})$, where $\widehat{\mathcal{T}}_{k,m}$ is the empirically adapted mean-field Bellman operator in Eq. 10.
Since the system is unknown, $\mathcal{T}_{k,m}$ and $\widehat{\mathcal{T}}_{k,m}$ cannot compute the direct expectation from the Bellman operator and instead draw $m$ random samples $s_{g}^{j}\!\sim\!P_{g}(\cdot|s_{g},a_{g})$ and $s_{i}^{j}\!\sim\!P_{l}(\cdot|s_{i},s_{g},a_{i})$ for each $j\!\in\![m]$, $i\!\in\!\Delta$:
$\begin{split}&{\widetilde{\mathcal{T}}}_{k,m}\widehat{Q}_{k,m}^{t}(s_{g},{s_{\Delta}},a_{g},a_{\Delta})=r_{\Delta}(s,a)+\frac{\gamma}{m}\sum_{j\in[m]}\max_{\begin{subarray}{c}a_{g}^{\prime}\in\mathcal{A}_{g},a_{\Delta}^{\prime}\in\mathcal{A}_{l}^{k}\end{subarray}}\widehat{Q}_{k,m}^{t}(s_{g}^{j},{s_{\Delta}^{j}},a_{g}^{\prime},a_{\Delta}^{\prime}).\end{split}$ (9)
$\begin{split}&\widehat{\mathcal{T}}_{k,m}\widehat{Q}_{k,m}^{t}(s_{g},s_{1},F_{z_{\widetilde{\Delta}}},a_{1},a_{g})=r_{\Delta}(s,a)+\frac{\gamma}{m}\sum_{j\in[m]}\max_{a_{g}^{\prime}\in\mathcal{A}_{g},a_{1}^{\prime}\in\mathcal{A}_{l},F_{a_{\widetilde{\Delta}}^{\prime}}\in\mu_{k-1}(\mathcal{A}_{l})}\widehat{Q}_{k,m}^{t}(s_{g}^{j},s_{1}^{j},F_{s_{\widetilde{\Delta}}^{j},a_{\widetilde{\Delta}^{\prime}}},a_{1}^{\prime},a_{g}^{\prime}).\end{split}$ (10)
$\widehat{Q}_{k,m}^{t}$ depends on $s_{\Delta}$ and $a_{\Delta}$ through $F_{z_{\Delta}}$, and $\widetilde{\mathcal{T}}_{k,m}$/$\widehat{\mathcal{T}}_{k,m}$ are $\gamma$-contractive (Lemma B.8). So, Algorithms 1 and 2 apply value iteration with their Bellman operator until $\widehat{Q}_{k,m}$ converges to a fixed point satisfying $\widetilde{\mathcal{T}}_{k,m}\widehat{Q}_{k,m}^{\mathrm{est}}\!=\!\widehat{Q}_{k,m}^{\mathrm{est}}$ and $\widehat{\mathcal{T}}_{k,m}\widehat{Q}_{k,m}^{\mathrm{est}}\!=\!\widehat{Q}_{k,m}^{\mathrm{est}}$, yielding equivalent deterministic policies $\widehat{\pi}^{\mathrm{est}}_{k,m}(s_{g},{s_{\Delta}})$ and $\widehat{\pi}^{\mathrm{est}}_{k,m}(s_{g},s_{1},F_{s_{\widetilde{\Delta}}})$:
$\widehat{\pi}^{\mathrm{est}}_{k,m}(s_{g},{s_{\Delta}})=\arg\max_{a_{g}\in\mathcal{A}_{g},a_{\Delta}\in\mathcal{A}_{l}^{k}}\widehat{Q}^{\mathrm{est}}_{k,m}(s_{g},{s_{\Delta}},a_{g},a_{\Delta})$ (11)
$\widehat{\pi}^{\mathrm{est}}_{k,m}(s_{g},s_{1},F_{s_{\widetilde{\Delta}}})=\arg\max_{a_{g}\in\mathcal{A}_{g},a_{1}\in\mathcal{A}_{l},F_{a_{\widetilde{\Delta}}}\in\mu_{k-1}(\mathcal{A}_{l})}\widehat{Q}^{\mathrm{est}}_{k,m}(s_{g},s_{1},F_{z_{\widetilde{\Delta}}},a_{1},a_{g})$ (12)
@@ 3 Method and Theoretical Results
Algorithm 3 (Online implementation). Here, Algorithm 3 (SUBSAMPLE-MFQ: Execution) randomly samples $\Delta\!\sim\!\mathcal{U}\binom{[n]}{k}$ at each time step and uses action $a_{g}\!\sim\!\widehat{\pi}_{k,m}^{\mathrm{est}}(s_{g},F_{s_{\Delta}})$ to get reward $r(s,a_{g})$. This procedure of first sampling $\Delta$ and then applying $\widehat{\pi}^{\mathrm{est}}_{k,m}$ is denoted by a stochastic policy ${\pi}_{k,m}^{\mathrm{est}}(a|s)=[{\pi}_{k,m}^{\mathrm{est}}(a_{g}|s),{\pi}_{k,m}^{\mathrm{est}}(a_{l}|s)]$, where ${\pi}_{k,m}^{\mathrm{est}}(a_{g}|s)$ is the global agent’s action distribution and ${\pi}_{k,m}^{\mathrm{est}}(a_{l}|s)$ is the local agent’s action distribution:
${\pi}_{k,m}^{\mathrm{est}}(a_{g}|s)=\frac{1}{\binom{n}{k}}\sum_{\Delta\in\binom{[n]}{k}}\mathbbm{1}(\widehat{\pi}_{k,m}^{\mathrm{est}}(s_{g},s_{\Delta})=a)$ (13)
${\pi}_{k,m}^{\mathrm{est}}(a_{i}|s)=\frac{1}{\binom{n-1}{k-1}}\sum_{\widetilde{\Delta}\in\binom{[n]\setminus i}{k-1}}\mathbbm{1}(\widehat{\pi}_{k,m}^{\mathrm{est}}(s_{g},s_{i},F_{s_{\widetilde{\Delta}}})=a_{i})$ (14)
Then, each agent transitions to their next state based on Eq. 1.
Algorithm 1 SUB-SAMPLE-MFQ: Learning (if $|\mathcal{Z}_{l}|^{k-1}\leq k^{|\mathcal{Z}_{l}|}$)
0:  A multi-agent system as described in Section 2. Parameter $T$ for the number of iterations in the initial value iteration step. Sampling parameters $k\in[n]$ and $m\in\mathbb{N}$. Discount parameter $\gamma\in(0,1)$. Oracle $\mathcal{O}$ to sample $s_{g}^{\prime}\sim{P}_{g}(\cdot|s_{g},a_{g})$ and $s_{i}^{\prime}\sim{P}_{l}(\cdot|s_{i},s_{g},a_{i})$ for all $i\in[n]$.
1:  Uniformly sample $\Delta\subseteq[n]$ such that $|\Delta|=k$.
2:  Initialize $\widehat{Q}^{0}_{k,m}(s_{g},s_{\Delta},a_{g},a_{\Delta})=0$ for $(s_{g},s_{\Delta},a_{g},a_{\Delta})\in\mathcal{S}_{g}\times\mathcal{S}_{l}^{k}\times\mathcal{A}_{g}\times\mathcal{A}_{l}^{k}$.
3:  for $t=1$ to $T$ do
4:     for $(s_{g},s_{\Delta},a_{g},a_{\Delta})\in\mathcal{S}_{g}\times\mathcal{S}_{l}^{k}\times\mathcal{A}_{g}\times\mathcal{A}_{l}^{k}$ do
5:        $\widehat{Q}^{t+1}_{k,m}(s_{g},s_{\Delta},a_{g},a_{\Delta})=\widetilde{\mathcal{T}}_{k,m}\widehat{Q}^{t}_{k,m}(s_{g},s_{\Delta},a_{g},a_{\Delta})$
6:  Return $\widehat{Q}_{k,m}^{T}$. For all $s_{g},s_{\Delta}\in\mathcal{S}_{g}\times\mathcal{S}_{l}^{k}$, let $\widehat{\pi}_{k,m}^{\mathrm{est}}(s_{g},s_{\Delta})=\arg\max_{a_{g}\in\mathcal{A}_{g},a_{\Delta}\in\mathcal{A}_{l}^{k}}\widehat{Q}_{k,m}^{T}(s_{g},s_{\Delta},a_{g},a_{\Delta})$.
Algorithm 2 SUBSAMPLE-MFQ: Learning (if $|\mathcal{Z}_{l}|^{k-1}<k^{|\mathcal{Z}_{l}|}$)
1:  Set $\widetilde{\Delta}=\{2,\dots,k\}$.
2:  Set $\mu_{k-1}(\mathcal{Z}_{l})=\{\frac{b}{k-1}:b\in\{0,1,\dots,k-1\}\}^{|\mathcal{S}_{l}|\times|\mathcal{A}_{l}|}$.
3:  Set $\widehat{Q}^{0}_{k,m}(s_{g},s_{1},F_{z_{\widetilde{\Delta}}},a_{1},a_{g})=0$, for $(s_{g},s_{1},F_{z_{\widetilde{\Delta}}},a_{1},a_{g})\in\mathcal{S}_{g}\times\mathcal{S}_{l}\times\mu_{k-1}(\mathcal{Z}_{l})\times\mathcal{A}_{l}\times\mathcal{A}_{g}$.
4:  for $t=1$ to $T$ do
5:     for $(s_{g},s_{1},F_{z_{\widetilde{\Delta}}},a_{1},a_{g})\in\mathcal{S}_{g}\times\mathcal{S}_{l}\times\mu_{k-1}(\mathcal{Z}_{l})\times\mathcal{A}_{l}\times\mathcal{A}_{g}$ do
6:        $\widehat{Q}^{t+1}_{k,m}(s_{g},s_{1},F_{z_{\widetilde{\Delta}}},a_{1},a_{g})=\widehat{\mathcal{T}}_{k,m}\widehat{Q}^{t}_{k,m}(s_{g},s_{1},F_{z_{\widetilde{\Delta}}},a_{1},a_{g})$
7:  $\forall(s_{g},s_{i},F_{s_{\widetilde{\Delta}}})\in\mathcal{S}_{g}\times\mathcal{S}_{l}\times\mu_{k-1}(\mathcal{S}_{l})$, let $\widehat{\pi}_{k,m}^{\mathrm{est}}(s_{g},s_{i},F_{s_{\widetilde{\Delta}}})=\operatorname*{arg\,max}_{a_{g}\in\mathcal{A}_{g},a_{i}\in\mathcal{A}_{l},F_{a_{\widetilde{\Delta}}}\in\mu_{k-1}{(\mathcal{A}_{l})}}\widehat{Q}_{k,m}^{T}(s_{g},s_{1},F_{z_{\widetilde{\Delta}}},a_{1},a_{g})$.
We define the greedy deterministic policy for the $k$-local-agent subsystem by: $\widehat{\pi}_{k,m}^{\mathrm{est}}(s_{g},s_{\Delta})$ by:
$\widehat{\pi}_{k,m}^{\mathrm{est}}(s_{g},s_{\Delta}):=\arg\max_{{a}_{g}^{*},{a}_{\Delta}^{*}}\widehat{Q}_{k,m}^{T}(s_{g},s_{\Delta},{a}_{g}^{*},{a}^{*}_{\Delta}):=([\widehat{\pi}_{k,m}^{\mathrm{est}}(s_{g},s_{\Delta})]_{g},[\widehat{\pi}_{k,m}^{\mathrm{est}}(s_{g},s_{i},s_{\Delta\setminus i})]_{l}),$ (15)
where $[\widehat{\pi}_{k,m}^{\mathrm{est}}(s_{g},s_{\Delta})]_{g}$ reads the maximizer $a_{g}^{*}$, and $[\widehat{\pi}_{k,m}^{\mathrm{est}}(s_{g},s_{i},s_{\Delta\setminus i})]_{l}$ reads the maximizer $a_{i}^{*}$. Then, for the $n$-agent system, the global agent samples local agents $\Delta$ uniformly from (${[n]\choose k}$ to derive action $a_{g}(t)=[\widehat{\pi}_{k,m}^{\mathrm{est}}(s_{g},s_{\Delta})]_{g}$, and allows agent $i$ to choose agents $\Delta_{i}$ uniformly from (${[n]\setminus i\choose k-1}$ to derive action $a_{i}(t)=[\widehat{\pi}_{k,m}^{\mathrm{est}}(s_{g},(s_{\Delta_{i}},s_{i})]_{l}$.
Algorithm 3 SUB-SAMPLE-MFQ: Execution
@@
0:  A multi-agent system as described in Section 2. Parameter $T^{\prime}$ for the number of iterations for the decision-making sequence. Hyperparameter $k\in[n]$. Discount parameter $\gamma$. Policy $\widehat{\pi}^{\mathrm{est}}_{k,m}(s_{g},F_{s_{\Delta}})$.
1:  Sample $(s_{g}(0),s_{[n]}(0))\sim s_{0}$, where $s_{0}$ is a distribution on the initial global state $(s_{g},s_{{[n]}})$
2:  Initialize the total reward $R_{0}=0$.
3:  Policy $\pi_{k}^{\mathrm{est}}(s)$ is defined as follows:
4:  for $t=0$ to $T^{\prime}$ do
5:     Choose $\Delta$ uniformly at random from (${[n]\choose k}$ and let $a_{g}(t)=[\widehat{\pi}_{k,m}^{\mathrm{est}}(s_{g}(t),s_{\Delta}(t))]_{g}$.
6:     for $i=1$ to $n$ do
7:        Choose $\Delta_{i}$ uniformly at random from (${[n]\setminus i\choose k-1}$ and let $a_{i}(t)=[\widehat{\pi}_{k,m}^{\mathrm{est}}(s_{g}(t),s_{i}(t),s_{\Delta_{i}}(t))]_{l}$.
8:     Let $s_{g}(t+1)\sim P_{g}(\cdot|s_{g}(t),a_{g}(t))$.
9:     Let $s_{i}(t+1)\sim P_{l}(\cdot|s_{i}(t),s_{g}(t),a_{i}(t))$, for all $i\in[n]$.
10:     $R_{t+1}=R_{t}+\gamma^{t}\cdot r(s,a)$
3.1 Theoretical Guarantee

### v3

@@ 3 Algorithmic Approach: Subsampled Value Iteration
Since mean-field $Q$-learning maintains the same updates as standard $Q$-learning (from Lemma C.19 which follows by noting that the $Q$-function is permutation-invariant in the homogeneous local agents), the deterministic policies $\hat{\pi}_{k,m}^{\mathrm{est}}(s_{g},s_{\Delta})$ and $\hat{\pi}_{k,m}^{\mathrm{est}}(s_{g},s_{1},F_{s_{\bar{\Delta}}})$ are equivalent.
@@
Algorithm 2 (Online implementation). In Algorithm 2, (SUBSAMPLE-MFQ: Execution) the global agent samples local agents $\Delta(t)\sim\mathcal{U}\binom{[n]}{k}$ at each step to derive action $a_{g}(t)=[\hat{\pi}_{k,m}^{\mathrm{est}}(s_{g},s_{\Delta}(t))]_{g}$, and each local agent $i$ samples other local agents $\Delta_{i}(t)\sim\mathcal{U}{[n]\setminus i\choose k-1}$ to derive action $a_{i}(t)=[\hat{\pi}_{k,m}^{\mathrm{est}}(s_{g},s_{i},s_{\Delta}(t))]_{l}$. The system then incurs a reward $r(s,a)$. This procedure of first sampling agents and then applying $\hat{\pi}^{\mathrm{est}}_{k,m}$ is denoted by a stochastic policy ${\pi}_{k,m}^{\mathrm{est}}(a|s)$, where ${\pi}_{k,m}^{\mathrm{est}}(a_{g}|s)$ is the global agent’s action distribution and ${\pi}_{k,m}^{\mathrm{est}}(a_{l}|s)$ is the local agent’s action distribution:
$\!\!{\pi}_{k,m}^{\mathrm{est}}(a_{g}|s)\!$ $=\!\frac{1}{\binom{n}{k}}\sum_{\Delta\in\binom{[n]}{k}}\mathbf{1}(\hat{\pi}_{k,m}^{\mathrm{est}}(s_{g},s_{\Delta})=a)$ (9)
$\!\!{\pi}_{k,m}^{\mathrm{est}}(a_{i}|s)\!$ $=\!\frac{1}{\binom{n-1}{k-1}}\!\!\sum_{\tilde{\Delta}\in\binom{[n]\setminus i}{k-1}}\!\!\!\!\mathbf{1}({\pi}_{k,m}^{\mathrm{est}}(s_{g},s_{i},F_{s_{\tilde{\Delta}}})\!=\!a_{i})\!$ (10)
The agents then transition to their next states.
Algorithm 2 SUBSAMPLE-MFQ: Execution
@@
0:  Parameter $T^{\prime}$ for the number of iterations for the decision-making sequence. Sampling parameter $k\in[n],m\in\mathbb{N}$. Discount factor $\gamma$. Policy $\hat{\pi}^{\mathrm{est}}_{k,m}(s_{g},F_{s_{\Delta}})$.
1:  Learn $\hat{\pi}_{k,m}^{\mathrm{est}}$ from Algorithm 1.
2:  Sample $(s_{g}(0),s_{[n]}(0))\sim s_{0}$, where $s_{0}$ is a distribution on the initial global state $(s_{g},s_{{[n]}})$
3:  Initialize the total reward $R_{0}=0$.
4:  Policy $\pi_{k,m}^{\mathrm{est}}(s)$ is defined as follows:
5:  for $t=0$ to $T^{\prime}$ do
6:     Choose $\Delta$ uniformly at random from (${[n]\choose k}$ and let $a_{g}(t)=[\hat{\pi}_{k,m}^{\mathrm{est}}(s_{g}(t),s_{\Delta}(t))]_{g}$.
7:     for $i=1$ to $n$ do
8:        Choose $\Delta_{i}$ uniformly at random from (${[n]\setminus i\choose k-1}$ and let $a_{i}(t)=[\hat{\pi}_{k,m}^{\mathrm{est}}(s_{g}(t),s_{i}(t),s_{\Delta_{i}}(t))]_{l}$.
9:     Let $s_{g}(t+1)\sim P_{g}(\cdot|s_{g}(t),a_{g}(t))$.
10:     Let $s_{i}(t+1)\sim P_{l}(\cdot|s_{i}(t),s_{g}(t),a_{i}(t))$, $\forall i\in[n]$.
11:     $R_{t+1}=R_{t}+\gamma^{t}\cdot r(s,a)$
4 Theoretical Guarantees and Analysis Approach